OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

perpendicular slop = 1/2 y-y1= m(x-x1) y-3= (1/2)(x-5) are they asking for slope intercept form/standard/or point slope form?

OpenStudy (anonymous):

He y sorry I couldn't connect so I couldn't reply or anything. I got it though thank you